OpenStudy (dragontears1128):

Please help and ill medal and fan Which of the following is in the solution set of y < 1x - 2 Select one: a. (9, -3) b. (12, 23) c. (11, 23) d. (9, 17)

OpenStudy (hawk):

hi

OpenStudy (hawk):

Idk answer

OpenStudy (jadeishere):

Did you just change the question? Or is my computer is glitching?

OpenStudy (dragontears1128):

i took one away

OpenStudy (jadeishere):

How did you do that without closing the question?

OpenStudy (jadeishere):

Nevermind! That's not important, I'll figure it out later. :) This is called a Linear Inequality and is fairly easy to get a handle on. Kay, a simple way to do this would be to substitute the values given into the inequality, to make a true statement.

jhonyy9 (jhonyy9):

y<x-2 Select one: a. (9, -3) b. (12, 23) c. (11, 23) d. (9, 17) a. for x=9 => y<9-2 so y<7 so y=-3 what is less than 7 so this is true b. for x=12 =>y<12-2 so y<10 but there is y=23 so this is false c. for x=11 =>y<11-2 so y<9 but there is y=23 so this is false d. for x=9 =>y<9-2 so y<7 but there is y=17 so this is false hope helped